OpenStudy (anonymous):

What is the total amount owed on a loan of $1000 at a 4% interest rate for 30 months? My answer is 2200

OpenStudy (koalagurl1):

I think it is 1200

OpenStudy (koalagurl1):

Are you adding the $1000?

OpenStudy (anonymous):

yeah, Im kinda confue @KoalaGurl1

OpenStudy (koalagurl1):

then you're right

OpenStudy (koalagurl1):

4% of 1000 is 40

OpenStudy (anonymous):

Just did the math......your right

OpenStudy (anonymous):

4% = 4 / 100 = 0.04 What is 4% of 1000 = 4% * 1000 = 0.04 * 1000 = 40

OpenStudy (koalagurl1):

so 40*30 is 1200

OpenStudy (koalagurl1):

then you add the $1000 to $1200 and you get $2200

OpenStudy (koalagurl1):

You are right.

OpenStudy (anonymous):

Calculate the simple interest on a loan of $15,000 at 3.5% for 5 years. my answer is 2625 @KoalaGurl1

OpenStudy (koalagurl1):

is it 3.5% a year?

OpenStudy (anonymous):

states for 5years

OpenStudy (koalagurl1):

huh...

OpenStudy (anonymous):

$15,000 at 3.5% for 5 years

OpenStudy (koalagurl1):

well if its per year then it is 2625 which is what you said

OpenStudy (koalagurl1):

if it is per month then it is 31500

OpenStudy (anonymous):

thanks

OpenStudy (koalagurl1):

No problem, glad I could help! :)